OpenStudy (anonymous):

how do I solve 8x + 4y = 16 on a TI 89?

Nnesha (nnesha):

you hve to solve for y

Nnesha (nnesha):

what's the statement ? do you need to graph it ? or find x-intercept .y-intercept ?

OpenStudy (anonymous):

I just need to solve 8x+4y=16 and x+y=2 solve the system but I want to know how to do it on my TI89

Nnesha (nnesha):

okay to put that into calculator (not online) you have to solve both equations for y

Nnesha (nnesha):

then press y= button type both equation Y_1= first equation y_2= 2nd equation

Nnesha (nnesha):

then hit the graph button

Nnesha (nnesha):

 when you press y= button u will see y_1= y_2= y_3= so you need to solve 8x+4y=16 for y